One endpoint of a segment is (12,-8). The midpoint is (3,18) Find the coordinates of the other endpoint.
Answer by stanbon(48568)

One endpoint of a segment is (12,-8). The midpoint is (3,18) Find the coordinates of the other endpoint.
-------------------------------------------------------
Let the endpoint be (x,y).
Then (x+12)/2 = 3
x+12=6
x=-6
And (y-8)/2 = 18
y-8=36
y=44
The endpoint is (-6,44)
Chers,
Stan H.